OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x - 7 y = - one halfx - 2 y = 2x - 2 y = - one halfx - 7

OpenStudy (anonymous):

@DollyAcquah can you help me with this

OpenStudy (texaschic101):

y = 2x + 2 --- in y = mx + b form, the number in the m position is the slope. A parallel line will have the same slope. The slope in this equation is 2. (5,3) x = 5 and y = 3 slope(m) = 2 now we sub that into the formula : y = mx + b 3 = 2(5) + b 3 = 10 + b 3 - 10 = b -7 = b your parallel equation is : y = 2x - 7 any questions ?

OpenStudy (anonymous):

no thank you so much.

OpenStudy (anonymous):

do you mind checking the ones i did?

OpenStudy (texaschic101):

or you could have used the point slope formula : y - y1 = m(x - x1) and solved it that way also

OpenStudy (texaschic101):

I will check your work

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = -one halfx + 3 y = -one halfx + 6 y = 2x + 3 I got A

OpenStudy (texaschic101):

slope(m) = -1/2 (6,3) x = 6 and y = 3 y = mx + b 3 = -1/2(6) + b 3 = -3 + b 3 + 3 = b 6 = b perpendicular line is : y = -1/2x + 6 your getting your parallel lines and you perpendicular lines mixed up

OpenStudy (texaschic101):

when you are looking for a perpendicular line, you want the negative reciprocal slope. y = 2x + 2...slope here is 2. But for a perpendicular line, you want the negative reciprocal. All that means is " flip " the slope and change the sign. slope of 2, or 2/1, " flipped" and sign changed = -1/2. That is the slope you want.

OpenStudy (anonymous):

so if i have a negative i change it to a positive

OpenStudy (texaschic101):

yes -2.....flip the slope and change the sign = 1/2 3 ...flip the slope and change the sign = -1/3 4/5...flip slope and change sign = -5/4 ONLY DO THIS IF YOU WANT A PERPENDICULAR LINE

OpenStudy (texaschic101):

because a parallel line has the same slope

OpenStudy (texaschic101):

understand ?

OpenStudy (anonymous):

a little

OpenStudy (texaschic101):

if you are looking for a parallel line, the slope will be the same slope as the equation given. if you want a perpendicular line, the slope is the negative reciprocal of the slope given...that is when you flip the slope and change the sign and use the flipped slope as your new slope

OpenStudy (anonymous):

okay i will apply that to the next problem

OpenStudy (texaschic101):

cool....if you need help, just tag me :)

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). stick on this

OpenStudy (anonymous):

Stuck * @texaschic101